# A complete characterization of testable hypotheses

### v1

@@ 2 Finitely additive measures and proofs of the main results
We regard $\mathcal{P}$} and $\mathcal{Q}$} as subsets of $L^{1}=L^{1}(\gamma)$ by identifying measures $\mu$ with their $\gamma$-densities $f_{\mu}=d\mu/d\gamma$. The total variation norm is then the $L^{1}$ norm of the density, $\|\mu\|_{\textnormal{TV}}=\|f_{\mu}\|_{L^{1}}$. Set $r=\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\operatorname{co}({\mathcal{Q}}))=\inf\frac{1}{2}\|f_{\mu}-f_{\nu}\|_{L^{1}}$, where the infimum extends over all $\mu\in{\mathcal{P}}$ and $\nu\in{\mathcal{Q}}$. Thus the convex set $\operatorname{co}({\mathcal{P}})-\operatorname{co}({\mathcal{Q}})$ is disjoint from $B_{2r}$, the open ball of radius $2r$ in $L^{1}$. The Hahn–Banach theorem (see, e.g., Schaefer and Wolff (1999, Theorem II.9.1)) then yields a unit-norm element $\psi\in L^{\infty}(\gamma)$ such that ${\mathbb{E}}_{\gamma}[\psi(f_{\mu}-f_{\nu})]\geq 2r$ for all $\mu\in\operatorname{co}({\mathcal{P}})$, $\nu\in\operatorname{co}({\mathcal{Q}})$. We fix a version of $\psi$ such that $|\psi|\leq 1$, and define the test $\phi=(1+\psi)/2\in\Phi$. It follows that ${\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi]\geq r$ for all $\mu,\nu$ in $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$, and then in $\operatorname{\overline{co}*}({\mathcal{P}})$ and $\operatorname{\overline{co}*}({\mathcal{Q}})$ by weak-* continuity. We conclude that $\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),\operatorname{\overline{co}*}({\mathcal{Q}}))\geq r$. On the other hand, the reverse inequality holds because $r$ is the distance between the two smaller sets $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$.
@@ 3 The case of a singleton $\mathcal{Q}$}
and call it the effective null hypothesis. No test can have nontrivial power against any distribution in $\mathcal{P}_{\textnormal{eff}}$. Indeed, as a corollary of their main theorem, we have the following:
@@
On the other hand, we also have the following immediate corollary of Theorem 1.5:
@@
The above theorem explains why, with their focus on singleton $\mathcal{Q}$}, Larsson et al. (2025) managed to avoid ba: the qualitative behavior of testability (i.e. the existence of a nontrivial test) does not require ba. Further, the quantitative statements of Larsson et al. (2025) in terms of infimum KL divergence from $\nu$ to $\mathcal{P}$} also avoided ba. However, even for a singleton $\nu$, the TV distances in Theorem 1.5 could not have used $\mathcal{P}_{\textnormal{eff}}$ instead of $\operatorname{\overline{co}*}({\mathcal{P}})$; we provide an example soon.
In order to prove the above theorem, we must characterize the relationship between $\mathcal{P}_{\textnormal{eff}}$ and $\operatorname{\overline{co}*}({\mathcal{P}})$. We do this in the following theorem, whose (rather lengthy) proof is presented in Section 5. It immediately implies Theorem 3.3.
@@
${\mathcal{P}}=\left\{\frac{1}{2}\delta_{x}+\frac{1}{2}\delta_{-x}\colon x\in(0,1)\right\},\quad{\mathcal{Q}}=\{\bar{\nu}\},\quad\bar{\nu}=\mathrm{Uni}[0,1].$
@@
Consider now any $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})\cap{\mathcal{M}}_{1}$. Its restriction to [1,0][-1,0] is a positive countably additive measure on $\textnormal{Pow}([-1,0])$, and is therefore concentrated on a countable set $S_{0}\subset[-1,0]$ thanks to Proposition A.3. The symmetry property applied with $A=-S_{0}$ then shows that $\mu$ is concentrated on $S=S_{0}\cup(-S_{0})$. But since any countable set is a $\bar{\nu}$-nullset, we have $\operatorname{d}_{\textnormal{TV}}(\mu,\bar{\nu})\geq\mu(S)-\bar{\nu}(S)=1$, and this yields the first part of (3.2).
We now focus on the second part of (3.2). The test $\phi=\bm{1}_{[0,1]}$ certifies that $\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),{\mathcal{Q}})$ is at least $\frac{1}{2}$, and the point is to show that this is achieved for some $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})$. To do this we start with a net $(\eta_{\alpha})$ of purely atomic measures on $(0,1)$ that converges to $\bar{\nu}$; this is possible thanks to Proposition A.1. The symmetrizations $\mu_{\alpha}(A\cup B)=\frac{1}{2}(\eta_{\alpha}(-A)+\eta_{\alpha}(B))$ for $A\in\textnormal{Pow}([-1,0])$ and $B\in\textnormal{Bor}([0,1])$ then belong to $\operatorname{co}({\mathcal{P}})$. On passing to a convergent subnet, we may assume that $(\mu_{\alpha})$ converges to some limit $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})$. The restriction of $\mu$ to $[0,1]$ is equal to $\bar{\nu}/2$ by construction, and it follows that $\operatorname{d}_{\textnormal{TV}}(\mu,\bar{\nu})=\frac{1}{2}$. This completes the proof of the second part of (3.2).
It is worth noting that the restriction of $\mu$ to [1,0][-1,0] must be purely finitely additive. Indeed, its countably additive part is atomless (due to the symmetry property and the fact that on $[0,1]$, $\mu$ is proportional to the uniform distribution) and therefore zero due to Proposition A.3.
@@
Example B.6 in Appendix B relies on more machinery than Example 3.5, but provides an even more striking example by constructing families $\mathcal{P}$} and $\mathcal{Q}$} such that
$\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}})\cap{\mathcal{M}}_{1},\operatorname{\overline{co}*}({\mathcal{Q}})\cap{\mathcal{M}}_{1})=1\quad\text{but}\quad\operatorname{\overline{co}*}({\mathcal{P}})\cap\operatorname{\overline{co}*}({\mathcal{Q}})\neq\emptyset.$
@@
4 Generalized tests
@@
It is useful to return to the basic example in the introduction, where $\Omega=[0,1]$, $\mathcal{F}$} is the Borel $\sigma$-algebra, ${\mathcal{P}}=\{\delta_{x}\colon x\in[0,1]\}$, and ${\mathcal{Q}}=\{\bar{\nu}\}$ with $\bar{\nu}$ the uniform distribution on $[0,1]$. The weak-* convex closure $\operatorname{\overline{co}*}({\mathcal{P}})$ contains all probability measures on $[0,1]$, in particular $\bar{\nu}$, so Theorem 1.5 implies that $R({\mathcal{P}},{\mathcal{Q}})=1$. Thus there is no nontrivial test for this pair of hypotheses. On the other hand, we have $\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\bar{\nu})=1$, so Le Cam’s Theorem 1.11 implies that there is a generalized test $\phi^{*}$ that perfectly separates $\mathcal{P}$} and $\mathcal{Q}$}.
In fact, it is not hard to construct such a generalized test. For any $\mu\in{\mathcal{M}}$, let $\mu^{a}$ be its absolutely continuous part with respect to $\bar{\nu}$. The map $\mu\mapsto\mu^{a}$ is linear and continuous (even a contraction) in the total variation norm. Now define $\phi^{*}(\mu)=\mu^{a}(\Omega)$, the total mass of $\mu^{a}$. Then $\phi^{*}$ is a continuous linear functional on $\mathcal{M}$}, and $\phi^{*}(\mu)\in[0,1]$ for all $\mu\in{\mathcal{M}}_{1}$. That is, $\phi^{*}$ is a generalized test. Furthermore, $\phi^{*}(\bar{\nu})=1$ and $\phi^{*}(\mu)=0$ for all $\mu\in{\mathcal{P}}$.
@@
5 Proof of Theorem 3.4
The first part of Theorem 3.4 states that every element of $\mathcal{P}_{\textnormal{eff}}$ is setwise dominated by some maximal element of $\mathcal{P}_{\textnormal{eff}}$. This is shown in Lemma 5.4 below, building on several auxiliary results which we now develop.
Lemma 5.1.
@@
By Lemma 5.1, $\mu_{\alpha}$ converges in the weak-* topology to some nonnegative countable additive measure $\mu$. This measure belongs to $\mathcal{P}_{\textnormal{eff}}$ because the property $\int fd\mu_{\alpha}\leq 1$ for all $f\in{\mathcal{E}}_{b}$ is preserved under weak-* convergence. ∎
@@
for all $n$. We now apply Lemma 5.2 to get that $\mu_{n}$ converges in the weak-* topology to some $\mu^{\prime}\in\mathcal{P}_{\textnormal{eff}}$ with $\mu^{\prime}\geq\mu_{n}$ for all $n$. Set
@@
Starting from $\mu_{0}$, we construct elements $\mu_{\alpha}\in\mathcal{P}_{\textnormal{eff}}$ for $\alpha\leq\gamma$ by transfinite recursion as follows. If $\alpha=\beta+1$ is a successor ordinal with $\alpha<\gamma$, we let $\mu_{\alpha}$ be an $A_{\alpha}$-maximal element of $\mathcal{P}_{\textnormal{eff}}$ such that $\mu_{\alpha}\geq\mu_{\beta}$. This is possible thanks to Lemma 5.3. If $\alpha=\beta+1=\gamma$, we simply set $\mu_{\alpha}=\mu_{\beta}$. If $\alpha$ is a limit ordinal, we first define $\nu_{\alpha}$ as the weak-* limit of the bounded increasing net $(\mu_{\beta})_{\beta<\alpha}\subset\mathcal{P}_{\textnormal{eff}}$, which exists and belongs to $\mathcal{P}_{\textnormal{eff}}$ thanks to Lemma 5.2. If $\alpha<\gamma$ we then get an $A_{\alpha}$-maximal $\mu_{\alpha}\in\mathcal{P}_{\textnormal{eff}}$ with $\mu_{\alpha}\geq\nu_{\alpha}$ from Lemma 5.3. If $\alpha=\gamma$, we simply set $\mu_{\alpha}=\nu_{\alpha}$. Now, for each $\alpha<\gamma$ we have $\mu_{\alpha}\leq\mu_{\gamma}$, so that $\mu_{\gamma}$ is also $A_{\alpha}$-maximal. Because $\{A_{\alpha}\colon\alpha<\gamma\}={\mathcal{F}}$, $\mu_{\gamma}$ is in fact $A$-maximal for every $A\in{\mathcal{F}}$. In other words, $\mu_{\gamma}$ is a maximal element of $\mathcal{P}_{\textnormal{eff}}$ that dominates $\mu_{0}$. ∎
The remaining parts of Theorem 3.4 require the introduction of a third object: one must view the null hypothesis $\mathcal{P}$} as a subset of ba and consider the following extension of the effective null,
$\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\{\mu\in\textnormal{{ba}}_{+}\colon\int fd\mu\leq 1\text{ for all }f\in{\mathcal{E}}_{b}\}.$ (5.1)
Lemma 5.5.
@@
${\mathcal{E}}_{b}=({\mathcal{P}}-\textnormal{{ba}}_{+})^{\circ}.$ (5.2)
The right-hand side equals $\{f\in{\mathcal{L}}\colon\int fd(\mu-\nu)\leq 1\text{ for all }\mu\in{\mathcal{P}},\nu\in\textnormal{{ba}}_{+}\}$, which certainly contains ${\mathcal{E}}_{b}$. Conversely, suppose $f$ belongs to the right-hand side. Taking $\nu=0$ we see that $\int fd\mu\leq 1$ for all $\mu\in{\mathcal{P}}$. Moreover, for any $\omega\in\Omega$ we may take an arbitrary $\mu\in{\mathcal{P}}$ and set $\nu=\mu+t\delta_{\omega}$ for $t>0$ to get $f(\omega)\geq-t^{-1}$. Sending $t$ to infinity shows that $f\in{\mathcal{L}}_{+}$, and hence that $f\in{\mathcal{E}}_{b}$. This establishes (5.2). Next, the definition (5.1) states that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\textnormal{{ba}}_{+}\cap{\mathcal{E}}_{b}^{\circ}$. Together with (5.2), this yields
@@
$\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}}).$ (5.3)
@@
We prove (5.3); the remaining statement is then immediate. In view of Lemma 5.5, it suffices to show that
$\textnormal{{ba}}_{+}\cap\operatorname{\overline{co}*}({\mathcal{P}}-\textnormal{{ba}}_{+})\subset\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}})\subset\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}.$ (5.4)
@@
We now prove the second inclusion in (5.4). It is clear from the definition (5.1) that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}$ is convex, weak-* closed, and contains $\mathcal{P}$}. Therefore it also contains $\operatorname{\overline{co}*}({\mathcal{P}})$. It is also clear that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}$ is solid, so it must contain $\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}})$ as well. This completes the proof of (5.4), and hence of (5.3). ∎
We now turn to the second part of Theorem 3.4; this is content of the following lemma. This also yields the final sentence of the theorem.
Lemma 5.7.

### v2

@@ 2 Finitely additive measures and proofs of the main results
We regard $\mathcal{P}$} and $\mathcal{Q}$} as subsets of $L^{1}=L^{1}(\gamma)$ by identifying measures $\mu$ with their $\gamma$-densities $f_{\mu}=d\mu/d\gamma$. The total-variation norm is then the $L^{1}$ norm of the density, $\|\mu\|_{\textnormal{TV}}=\|f_{\mu}\|_{L^{1}}$. Set $r=\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\operatorname{co}({\mathcal{Q}}))=\inf\frac{1}{2}\|f_{\mu}-f_{\nu}\|_{L^{1}}$, where the infimum extends over all $\mu\in{\mathcal{P}}$ and $\nu\in{\mathcal{Q}}$. Thus the convex set $\operatorname{co}({\mathcal{P}})-\operatorname{co}({\mathcal{Q}})$ is disjoint from $B_{2r}$, the open ball of radius $2r$ in $L^{1}$. The Hahn–Banach theorem (see, e.g., Schaefer and Wolff (1999, Theorem II.9.1)) then yields a unit-norm element $\psi\in L^{\infty}(\gamma)$ such that ${\mathbb{E}}_{\gamma}[\psi(f_{\mu}-f_{\nu})]\geq 2r$ for all $\mu\in\operatorname{co}({\mathcal{P}})$, $\nu\in\operatorname{co}({\mathcal{Q}})$. We fix a version of $\psi$ such that $|\psi|\leq 1$, and define the test $\phi=(1+\psi)/2\in\Phi$. It follows that ${\mathbb{E}}_{\mu}[\phi]-{\mathbb{E}}_{\nu}[\phi]\geq r$ for all $\mu,\nu$ in $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$, and then in $\operatorname{\overline{co}*}({\mathcal{P}})$ and $\operatorname{\overline{co}*}({\mathcal{Q}})$ by weak-* continuity. We conclude that $\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),\operatorname{\overline{co}*}({\mathcal{Q}}))\geq r$. On the other hand, the reverse inequality holds because $r$ is the distance between the two smaller sets $\operatorname{co}({\mathcal{P}})$ and $\operatorname{co}({\mathcal{Q}})$.
@@ 3 The case of a singleton $\mathcal{Q}$}
and call it the effective null hypothesis. No test can have nontrivial power against any distribution in $\mathcal{P}_{\textnormal{eff}}$. Indeed, as a corollary of their main theorem, we have the following.
@@
On the other hand, we also have the following immediate corollary of Theorem 1.5.
@@
The above theorem explains why, with their focus on a singleton ${\mathcal{Q}}=\{\nu\}$, Larsson et al. (2025) managed to avoid ba: the qualitative behavior of testability (i.e. the existence of a nontrivial test) does not require ba. Further, the quantitative statements of Larsson et al. (2025) in terms of infimum KL divergence from $\nu$ to $\mathcal{P}$} also avoided ba. However, even for a singleton ${\mathcal{Q}}=\{\nu\}$, the TV distances in Theorem 1.5 could not have used $\mathcal{P}_{\textnormal{eff}}$ instead of $\operatorname{\overline{co}*}({\mathcal{P}})$; we provide an example soon.
In order to prove the above theorem, we must characterize the relationship between $\mathcal{P}_{\textnormal{eff}}$ and $\operatorname{\overline{co}*}({\mathcal{P}})$. We do this in the following theorem, whose (rather lengthy) proof is presented in Appendix D. It immediately implies Theorem 3.3.
@@
${\mathcal{P}}=\left\{\frac{1}{2}\delta_{x}+\frac{1}{2}\delta_{-x}\colon x\in(0,1)\right\},\quad{\mathcal{Q}}=\{\nu\},\quad\nu=\mathrm{Uni}[0,1].$
@@
Consider now any $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})\cap{\mathcal{M}}_{1}$. Its restriction to [1,0][-1,0] is a positive countably additive measure on $\textnormal{Pow}([-1,0])$, and is therefore concentrated on a countable set $S_{0}\subset[-1,0]$ thanks to Proposition C.3. The symmetry property applied with $A=-S_{0}$ then shows that $\mu$ is concentrated on $S=S_{0}\cup(-S_{0})$. But since any countable set is a $\nu$-nullset, we have $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)\geq\mu(S)-\nu(S)=1$, and this yields the first part of (3.2).
We now focus on the second part of (3.2). The test $\phi=\bm{1}_{[0,1]}$ certifies that $\operatorname{d}_{\textnormal{TV}}(\operatorname{\overline{co}*}({\mathcal{P}}),{\mathcal{Q}})$ is at least $\frac{1}{2}$, and the point is to show that this is achieved for some $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})$. To do this we start with a net $(\eta_{\alpha})$ of purely atomic measures on $(0,1)$ that converges to $\nu$; this is possible thanks to Proposition C.1. The symmetrizations $\mu_{\alpha}(A\cup B)=\frac{1}{2}(\eta_{\alpha}(-A)+\eta_{\alpha}(B))$ for $A\in\textnormal{Pow}([-1,0])$ and $B\in\textnormal{Bor}([0,1])$ then belong to $\operatorname{co}({\mathcal{P}})$. On passing to a convergent subnet, we may assume that $(\mu_{\alpha})$ converges to some limit $\mu\in\operatorname{\overline{co}*}({\mathcal{P}})$. The restriction of $\mu$ to $[0,1]$ is equal to $\nu/2$ by construction, and it follows that $\operatorname{d}_{\textnormal{TV}}(\mu,\nu)=\frac{1}{2}$. This completes the proof of the second part of (3.2).
It is worth noting that the restriction of $\mu$ to [1,0][-1,0] must be purely finitely additive. Indeed, its countably additive part is atomless (due to the symmetry property and the fact that on $[0,1]$, $\mu$ is proportional to the uniform distribution) and therefore zero due to Proposition C.3.
@@
Acknowledgments
Appendix A Further comments on the choice of topology
In the introduction we showed why the TV topology and the weak convergence topology cannot, in general, be used to obtain characterizations of testability; see Examples 1.31.4 and the adjacent discussion. An alternative approach would be to look for topologies on the set $\Phi$ of all tests, making this set compact. This could then be leveraged to apply the minimax theorem without compactness of the closed convex hulls of $\mathcal{P}$} and $\mathcal{Q}$}. This idea works in the dominated setting of Theorem 1.1, and Le Cam’s approach to embed $\Phi$ into the larger, compact, space of generalized tests can be viewed as an extension of this idea to the general case, with important drawbacks as discussed in Appendix B.
There are, however, topologies that make $\Phi$ itself compact in the general non-dominated setting. One prominent example is the topology of pointwise convergence. This is the weakest topology that makes the evaluation maps $\phi\mapsto\phi(\omega)$ continuous for every $\omega\in\Omega$. Equivalently, this is the product topology on $0,1]^{\Omega}$, which is compact by Tychonoff’s theorem. The trouble with this topology is that the linear functions $\phi\mapsto{\mathbb{E}}_{\mu}[\phi]$ fail to be continuous (or semicontinuous) in general.
For example, take $\Omega=[0,1]$ and $\mu$ Lebesgue measure. For every finite subset $A=\{\omega_{1},\ldots,\omega_{n}\}\subset[0,1]$, let $\phi_{A}=\bm{1}_{A}$ be the indicator function of $A$, which is a test. These tests form a net that converges to the constant test $\phi=1$ in the topology of pointwise convergence. However, ${\mathbb{E}}_{\mu}[\phi_{A}]=0$ for all $A$, while ${\mathbb{E}}_{\mu}[\phi]=1$, showing that the corresponding expectations do not converge to the expectation of the limiting test. Thus, even though $\Phi$ is compact, the semicontinuity hypothesis of the minimax theorem is not satisfied, and the theorem cannot be used.
To restore (semi-)continuity, one has to work with a stronger topology. A very strong topology is the one induced by the supremum norm $\|\phi\|_{\infty}=\sup_{\omega\in\Omega}|\phi(\omega)|$. The above linear maps $\phi\mapsto{\mathbb{E}}_{\mu}[\phi]$ are now continuous thanks to the bounded convergence theorem. However, it is now compactness of $\Phi$ that fails, again rendering the minimax theorem inapplicable. Indeed, given any sequence $\omega_{n}$ of mutually distinct points, the tests $\phi_{n}=\bm{1}_{\{\omega_{n}\}}$ are at a distance $\|\phi_{n}-\phi_{m}\|_{\infty}=1$ apart, and thus form a sequence with no convergent subsequence.
At this point it may be tempting to consider the topology $\sigma(\Phi,{\mathcal{M}})$, which is the weakest topology on $\Phi$ that makes the linear maps $\phi\mapsto{\mathbb{E}}_{\mu}[\phi]$ continuous for all $\mu\in{\mathcal{M}}$. Unfortunately, compactness of $\Phi$ still fails, and the net $\phi_{A}=\bm{1}_{A}$ indexed by finite subsets $A\subset\Omega=[0,1]$ again furnishes an example. Indeed, suppose for contradiction that compactness holds. Then there is a convergent subnet $\phi_{B}$ with some limit $\phi$. This limit is the constant function $\phi=1$. To see this, note that for each $\omega\in[0,1]$, $\phi_{B}(\omega)=1$ eventually, so $\phi(\omega)={\mathbb{E}}_{\delta_{\omega}}[\phi]=\lim_{B}{\mathbb{E}}_{\delta_{\omega}}[\phi_{B}]=1$, where we used that $\lim_{B}{\mathbb{E}}_{\mu}[\phi_{B}]={\mathbb{E}}_{\mu}[\phi]$ for every $\mu\in{\mathcal{M}}$ by definition of the topology. However, if we now take $\mu$ to be Lebesgue measure, we obtain the contradiction $0=\lim_{B}{\mathbb{E}}_{\mu}[\phi_{B}]={\mathbb{E}}_{\mu}[\phi]=1$. This shows that $\phi_{A}$ has no convergent subnet, and hence that $\Phi$ is not compact.
One thus arrives at the conclusion that either $\Phi$ has to be “compactified” (for example by enlarging it to generalized tests as proposed by Le Cam, with its drawbacks discussed in Appendix B), or one has to find a way to compactify the space of measures rather than tests. The latter is the route taken in the present paper (by enlarging the space of measures), which leads us naturally — even necessarily — to working with finitely additive measures.
We hope that this discussion gives a further perspective on the delicate interplay between the choice of ambient space, its topology, and how these must be tailored to fit the confines of the minimax theorem. Unfortunately, we are aware of several published works which incorrectly use minimax theorems to achieve important conclusions. For example Gül and Zoubir (2017) mistakenly invoke Sion’s minimax theorem after using the product topology on $0,1]^{\Omega}$, which as we discussed above destroys continuity. As another example Levy (2008, Chapter 6) invokes von Neumann’s minimax theorem after endowing $\Phi$ with $\|\cdot\|_{\infty}$, which as we discussed above renders $\Phi$ non-compact. While we know of several more examples, it is not our goal to point out all such errors in other works, but merely to point to the subtleties involved with the careful choices necessary for everything to work out. It would also be of interest to see whether the corresponding conclusions in these works are incorrect, or if their proofs can be fixed using techniques developed in this paper.
Appendix B Generalized tests
@@
It is useful to return to the basic example in the introduction, where $\Omega=[0,1]$, $\mathcal{F}$} is the Borel $\sigma$-algebra, ${\mathcal{P}}=\{\delta_{x}\colon x\in[0,1]\}$, and ${\mathcal{Q}}=\{\nu\}$ with $\nu=\mathrm{Uni}[0,1]$, the uniform distribution on $[0,1]$. The weak-* convex closure $\operatorname{\overline{co}*}({\mathcal{P}})$ contains all probability measures on $[0,1]$, in particular $\nu$, so Theorem 1.5 implies that $R({\mathcal{P}},{\mathcal{Q}})=1$. Thus there is no nontrivial test for this pair of hypotheses. On the other hand, we have $\operatorname{d}_{\textnormal{TV}}(\operatorname{co}({\mathcal{P}}),\nu)=1$, so Le Cam’s Theorem 1.11 implies that there is a generalized test $\phi^{*}$ that perfectly separates $\mathcal{P}$} and $\mathcal{Q}$}.
In fact, it is not hard to construct such a generalized test. For any $\mu\in{\mathcal{M}}$, let $\mu^{a}$ be its absolutely continuous part with respect to $\nu$. The map $\mu\mapsto\mu^{a}$ is linear and continuous (even a contraction) in the total-variation norm. Now define $\phi^{*}(\mu)=\mu^{a}(\Omega)$, the total mass of $\mu^{a}$. Then $\phi^{*}$ is a continuous linear functional on $\mathcal{M}$}, and $\phi^{*}(\mu)\in[0,1]$ for all $\mu\in{\mathcal{M}}_{1}$. That is, $\phi^{*}$ is a generalized test. Furthermore, $\phi^{*}(\nu)=1$ and $\phi^{*}(\mu)=0$ for all $\mu\in{\mathcal{P}}$.
@@
Appendix C Some measure-theoretic results
Proposition C.1.
Remark C.2.
Proof of Proposition C.1.
Proposition C.3.
$\mu(A)=\sum_{x\in A\cap S}\mu(\{x\}).$
Proof.
Appendix D Proof of Theorem 3.4
The first part of Theorem 3.4 states that every element of $\mathcal{P}_{\textnormal{eff}}$ is setwise dominated by some maximal element of $\mathcal{P}_{\textnormal{eff}}$. This is shown in Lemma D.4 below, building on several auxiliary results which we now develop.
Lemma D.1.
@@ Appendix C Some measure-theoretic results
By Lemma D.1, $\mu_{\alpha}$ converges in the weak-* topology to some nonnegative countable additive measure $\mu$. This measure belongs to $\mathcal{P}_{\textnormal{eff}}$ because the property $\int fd\mu_{\alpha}\leq 1$ for all $f\in{\mathcal{E}}_{b}$ is preserved under weak-* convergence. ∎
@@
for all $n$. We now apply Lemma D.2 to get that $\mu_{n}$ converges in the weak-* topology to some $\mu^{\prime}\in\mathcal{P}_{\textnormal{eff}}$ with $\mu^{\prime}\geq\mu_{n}$ for all $n$. Set
@@
Starting from $\mu_{0}$, we construct elements $\mu_{\alpha}\in\mathcal{P}_{\textnormal{eff}}$ for $\alpha\leq\gamma$ by transfinite recursion as follows. If $\alpha=\beta+1$ is a successor ordinal with $\alpha<\gamma$, we let $\mu_{\alpha}$ be an $A_{\alpha}$-maximal element of $\mathcal{P}_{\textnormal{eff}}$ such that $\mu_{\alpha}\geq\mu_{\beta}$. This is possible thanks to Lemma D.3. If $\alpha=\beta+1=\gamma$, we simply set $\mu_{\alpha}=\mu_{\beta}$. If $\alpha$ is a limit ordinal, we first define $\nu_{\alpha}$ as the weak-* limit of the bounded increasing net $(\mu_{\beta})_{\beta<\alpha}\subset\mathcal{P}_{\textnormal{eff}}$, which exists and belongs to $\mathcal{P}_{\textnormal{eff}}$ thanks to Lemma D.2. If $\alpha<\gamma$ we then get an $A_{\alpha}$-maximal $\mu_{\alpha}\in\mathcal{P}_{\textnormal{eff}}$ with $\mu_{\alpha}\geq\nu_{\alpha}$ from Lemma D.3. If $\alpha=\gamma$, we simply set $\mu_{\alpha}=\nu_{\alpha}$. Now, for each $\alpha<\gamma$ we have $\mu_{\alpha}\leq\mu_{\gamma}$, so that $\mu_{\gamma}$ is also $A_{\alpha}$-maximal. Because $\{A_{\alpha}\colon\alpha<\gamma\}={\mathcal{F}}$, $\mu_{\gamma}$ is in fact $A$-maximal for every $A\in{\mathcal{F}}$. In other words, $\mu_{\gamma}$ is a maximal element of $\mathcal{P}_{\textnormal{eff}}$ that dominates $\mu_{0}$. ∎
The remaining parts of Theorem 3.4 require the introduction of a third object: one must view the null hypothesis $\mathcal{P}$} as a subset of ba and consider an appropriate extension of the effective null, namely,
$\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\{\mu\in\textnormal{{ba}}_{+}\colon\int fd\mu\leq 1\text{ for all }f\in{\mathcal{E}}_{b}\}.$ (D.1)
Lemma D.5.
@@
${\mathcal{E}}_{b}=({\mathcal{P}}-\textnormal{{ba}}_{+})^{\circ}.$ (D.2)
The right-hand side equals $\{f\in{\mathcal{L}}\colon\int fd(\mu-\nu)\leq 1\text{ for all }\mu\in{\mathcal{P}},\nu\in\textnormal{{ba}}_{+}\}$, which certainly contains ${\mathcal{E}}_{b}$. Conversely, suppose $f$ belongs to the right-hand side. Taking $\nu=0$ we see that $\int fd\mu\leq 1$ for all $\mu\in{\mathcal{P}}$. Moreover, for any $\omega\in\Omega$ we may take an arbitrary $\mu\in{\mathcal{P}}$ and set $\nu=\mu+t\delta_{\omega}$ for $t>0$ to get $f(\omega)\geq-t^{-1}$. Sending $t$ to infinity shows that $f\in{\mathcal{L}}_{+}$, and hence that $f\in{\mathcal{E}}_{b}$. This establishes (D.2). Next, the definition (D.1) states that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\textnormal{{ba}}_{+}\cap{\mathcal{E}}_{b}^{\circ}$. Together with (D.2), this yields
@@
$\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}=\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}}).$ (D.3)
@@
We prove (D.3); the remaining statement is then immediate. In view of Lemma D.5, it suffices to show that
$\textnormal{{ba}}_{+}\cap\operatorname{\overline{co}*}({\mathcal{P}}-\textnormal{{ba}}_{+})\subset\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}})\subset\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}.$ (D.4)
@@
We now prove the second inclusion in (D.4). It is clear from the definition (D.1) that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}$ is convex, weak-* closed, and contains $\mathcal{P}$}. Therefore it also contains $\operatorname{\overline{co}*}({\mathcal{P}})$. It is also clear that $\mathcal{P}_{\textnormal{eff}}^{\textnormal{{ba}}}$ is solid, so it must contain $\operatorname{solid}\operatorname{\overline{co}*}({\mathcal{P}})$ as well. This completes the proof of (D.4), and hence of (D.3). ∎
We now turn to the second part of Theorem 3.4; this is content of the following lemma.
Lemma D.7.